HW 2.1 Graphical Summaries for Qualitative Data Previous 3 Next Question 5 of 5 (1 point) View problem in a pop-up 2.1 Section Exercise 27bc In a survey commissioned by an electronies company, U.S. computer users were asked how often E they backed up their computers hard drive. The following frequency distribution presents the results. Response More than once per month Once every 1-3 months Once every 4-6 months Once every 7-11 monts Once per year or less Never Frequency 392 421 293 80 369 798 Sh Ask Link Fc Download data Save Part 1 Which of the visual displays below is a relative frequeney bar graph? Submit A
Ask Cuba Iran Link The relative frequency bar graph is c Fa Part 2 out of 3 Construct a relative frequeney distribution. Round the answers to three decimal places. Save Response Relative Frequency Submit A More than once per month Once every 1-3 months Once every 4-6 months Once every 7-11 months Once per year or less Never 8 F3
Answer #1

Answer) Relative frequency is given by the frequency divided by the total sum of all the frequencies.

So, therefore,

the sum of all the frequencies = 392+421+293+80+369+798 = 2353

Relative frequency of more than once per month = 392/2353 = 0.166

Relative frequency of once every 1-3 months= 421/2353 = 0.178

Relative frequency of once every 4-6 months = 293/2353 = 0.124

Relative frequency of once 7-11 months = 80/2353 = 0.033

Relative frequency of once per year or less = 369/2353 = 0.156

Relative frequency of never = 798/2353 = 0.339